# Collaborative Deterministic–Probabilistic Forecasting for Real-World Spatiotemporal Systems

### v4

@@ 4.3 Residual Learning via Diffusion Model
The residual distribution of spatiotemporal data is not independently and identically distributed (i.i.d.) nor does it follow a fixed distribution, such as $\mathcal{N}(0,\mathbf{\sigma})$. Instead, it often exhibits complex spatiotemporal dependence and heterogeneity. We use the diffusion model to focus on learning the distribution of residual $\mathbf{r}^{ta}=\mathbf{x}^{ta}-\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}].$
Accordingly, the target data $\mathbf{x}^{ta}$ for diffusion models in Eqs. (1), (2), and (3) is replaced by $\mathbf{r}^{ta}$. We incorporate timestamp information as a condition in the denoising process and concatenate the context data $\mathbf{x}^{co}_{0}$ with noised residual $\mathbf{r}^{ta}_{n}$ as input to capture real-time fluctuations. Notably, no noise is added to $\mathbf{x}^{co}_{0}$ during diffusion training or inference. To model the spatial patterns of the residuals, we propose a scale-aware diffusion process to further distinguish the heterogeneity for different spatial units. In this section, we detail the calculation of $\mathbf{Q}$ and how it is integrated into the scale-aware diffusion process.
@@
(ii) Scale-aware diffusion process. The vanilla diffusion models assume a shared prior distribution $\mathcal{N}(0,I)$ across all spatial locations, failing to capture spatial heterogeneity. To further model such differences, we adopt the technique proposed by [21] to make the residual learning location-specific conditioned on ${\mathbf{Q}}$. Specifically, we redefine the noise distribution at the endpoint of the diffusion process as follows:
@@ 4.4 Training and Inference
Our training follows a two-stage training procedure: we first pretrain a deterministic model to predict the conditional mean, then train a diffusion model to capture the residual distribution. The full procedure is outlined in Algorithm 1. The inference consists of two paths: the pretrained deterministic model predicts the conditional mean, and the diffusion model estimates the residuals. Their outputs are combined to form the final prediction, as detailed in Algorithm 2.
@@ 5 Experiments
Baselines. We compare against six representative state-of-the-art baselines commonly adopted in spatiotemporal modeling, including: D3VAE [30], DiffSTG [57], TimeGrad [46], CSDI [53], DYffusion [47], and NPDiff [51]. Detailed descriptions of each baseline are provided in Appendix C.2.
Table 1: Short-term forecasting results in terms of CRPS, QICE, and IS. Bold indicates the best performance, while underlining denotes the second-best. DYffusion is limited to grid-format data, and ‘-’ denotes results that are not applicable.
@@
Metrics. To evaluate the performance, we employ two deterministic metrics, MAE and RMSE, along with three probabilistic metrics: CRPS, QICE and IS. For QICE, we set the number of QIs, denoted as $M_{\text{QIs}}$, to 10. We choose 10 bins for QICE to align with the original proposal [21]. which offers a balanced trade-off between granularity and stability. For IS, we choose a confidence level of 90% (i.e., $\alpha_{CI}=0.1$) following common practice in spatiotemporal forecasting tasks [53, 46].
Experimental configuration. We define the short-term forecasting task as predicting the next 12 time steps based on the previous 12 observations, following [51, 57]. Since the temporal granularity varies across datasets, the actual time duration corresponding to these 12 steps differs accordingly. In addition to the standard 12-step setting commonly used in spatiotemporal forecasting, we evaluate long-term forecasting by predicting 64 future steps based on the preceding 64 observations, following [63, 27, 40]. Detailed training and model configurations are provided in Appendix C.3.
@@ 5.1 Spatiotemporal Probabilistic Forecasting
Short-term forecasting. Table 1 presents the results of probabilistic metrics for selected datasets. Due to space constraints, the remaining results are in Appendix Table 6. As shown in Table 1, CoST consistently outperforms baseline methods across all evaluated datasets. Compared to the best-performing baseline methods on each dataset, CoST demonstrates an average improvement of 17.4% in CRPS and 46.6% in QICE metrics, indicating its superior ability to accurately capture the true distribution characteristics. Moreover, CoST achieves a 16.5% improvement in the IS metric, suggesting that its prediction intervals not only maintain compactness but also exhibit higher coverage, thereby better reflecting the uncertainty of data. Although certain individual metrics may not reach the optimal level on specific datasets, CoST consistently maintains performance comparable to the best methods. Beyond probabilistic metrics, we also report deterministic evaluation results (MAE and RMSE) in Table 2 and Appendix Table 7. The results show that our method achieves an average reduction of 7% in MAE and 6.1% in RMSE datasets. This suggests that the integration with a strong conditional mean estimator enables CoST to better capture regular patterns compared to other probabilistic baseline models.
Long-term forecasting. As shown in Appendix Table  8, CoST achieves substantial improvements in long-term forecasting under probabilistic metrics, with an improvement of 15.0% and 70.4% in terms of CRPS and QICE. Despite adopting a simple MLP architecture, CoST achieves higher overall accuracy than CSDI, a Transformer-based model tailored for capturing long-range dependencies. Furthermore, it provides significantly better training efficiency and inference speed, as detailed in Section 5.4. In addition, CoST performs well on deterministic metrics (Appendix Table  9), achieving an average reduction of 9.0% in MAE and 11.0% in RMSE compared to the best-performing baseline.
Table 3: Performance of different deterministic backbone models within the CoST framework. ’Diffusion (w/o m)’ denotes the results obtained using a single diffusion model.
@@
Framework generalization. To demonstrate the generality of CoST, we instantiate it with four representative spatiotemporal forecasting models: STID [49], STNorm [14], ConvLSTM [52], and iTransformer [36]. These models cover a diverse set of deep learning architectures, including CNNs, RNNs, MLPs, and Transformers. As shown in Table 3, CoST consistently enhances the performance of these backbones by effectively integrating deterministic and probabilistic modeling. Compared to using a single diffusion model, CoST yields more accurate predictions and better-calibrated uncertainty estimates, validating the framework’s broad applicability and effectiveness.
@@
Case study of SST forecasting. To assess our model’s ability to quantify uncertainty under complex climate dynamics, we evaluate its performance in a key region for ENSO-related Sea Surface Temperature (SST) forecasting. As shown in Figure 3, our model produces high-fidelity SST forecasts that closely match ground truth across both warm pool and cold tongue regions. In addition to accurate mean predictions, it provides well-calibrated uncertainty estimates, revealing elevated variance in the central equatorial Pacific, especially near 0° latitude and 140°–130°W, where sharp thermocline gradients and nonlinear feedbacks make forecasting particularly challenging. These high-uncertainty areas align with known regions of model divergence in climate science [39, 6, 7], demonstrating that our method delivers both accurate predictions and geophysically consistent uncertainty estimates.
@@ 5.2 Ablation Study
We perform an ablation study to assess the contribution of each proposed module. Specifically, we construct three model variants by progressively removing key components: (w/o s): removes the scale-aware diffusion process; (w/o q) excludes the customized fluctuation scale as a prior; (w/o m) removes the conditional mean predictor, relying solely on the diffusion model. We conduct experiments on two datasets and visualize the results in terms of CRPS and IS metrics, as illustrated in Appendix Figure 8. Results show that the deterministic predictor notably improves performance by capturing regular spatiotemporal patterns, while also reducing the diffusion model’s complexity. Adding the customized fluctuation scale further enhances accuracy, indicating its utility in providing valuable fluctuation information across different spatial units. And the scale-aware diffusion process enables the diffusion model to better utilize this condition.
@@ 5.3 Qualitative Analysis
As shown in Figure 5, the ground truth exhibits clear spatiotemporal multi-modality. In Figure 5(a), three peaks likely correspond to different time points or varying states at the same time. CoST accurately captures all three peaks, while CSDI only fits two, showing CoST’s superior multi-modal modeling. In Figure 5(b), both models capture two peaks, but CoST aligns better with the peak spacing in the true distribution, reflecting stronger temporal sensitivity. These strengths arise from CoST’s hybrid design: the diffusion component models residual uncertainty to capture multi-modal traits, while the deterministic backbone learns regular trends. See Appendix C.5.1 for more analysis and results.
@@
Analysis of prediction quality. To intuitively demonstrate the effectiveness of our predictions, we visualize results on the CrowdBJ dataset in Figure 6, comparing our model with the best baseline, CSDI. As shown in Figures 6 (a, c, f), our model, aided by a deterministic backbone, better captures regular spatiotemporal patterns. Meanwhile, the diffusion module enhances uncertainty modeling by focusing on residuals, as reflected in Figures 6 (b, d, e). Beyond sample-level comparison, we evaluate prediction interval calibration via dynamic quantile error curves on CrowdBM and MobileSH (Figure 5). For each confidence level $\alpha$, we compute the corresponding quantile interval and its Prediction Interval Coverage Probability (PICP). Closer alignment with the diagonal (black dashed line) indicates better calibration. Our model consistently outperforms CSDI in this regard.
@@ 5.4 Computational Cost
We benchmark training and inference time (including 50 sampling iterations and pretraining for our mean predictor) on the MobileSH dataset. As shown in Appendix Table LABEL:tbl:traintime, our method achieves markedly higher efficiency than existing probabilistic models in both training and inference. This efficiency is particularly advantageous for real-world applications such as mobile traffic prediction. Notably, CSDI leverages the Transformer’s expressive power but incurs substantial computational cost, limiting its applicability in time-sensitive settings.
@@ 6 Conclusion
In this work, we highlight the importance of probabilistic forecasting for complex spatiotemporal systems and propose CoST, a collaborative framework that integrates deterministic and diffusion models. By decomposing data into a conditional mean and residual component, CoST bridges deterministic and probabilistic modeling, enabling accurate capture of both regular patterns and uncertainties. Extensive experiments on seven real-world datasets show that CoST consistently outperforms state-of-the-art methods with an average improvement of 25%. Our approach offers an effective solution for combining precise pattern learning with uncertainty modeling in spatiotemporal forecasting.
@@
Appendix A Background
A.1 Glossary
@@
$N$ The number of diffusion steps.
$H$ Height of the grid-based data
$W$ Width of the grid-based data
@@
$\mathcal{K}$ $K$ The set of indices for the selected FFT components
@@ A.2 Spatiotemporal Data
Appendix B Methodology
B.1 Algorithm
@@
Appendix C Experiments
C.1 Datasets
Table 5: The basic information of grid-based spatio-temporal data.
@@ C.1 Datasets
Los-Speed Los Angeles Traffic speed 2012/03/01 - 2012/03/07 $207$ 5 minutes
@@ C.2 Baselines
  • D3VAE [30]: Aims at short-period and noisy time series forecasting. It combines generative modeling with a bidirectional variational auto-encoder, integrating diffusion, denoising, and disentanglement.
@@
  • DiffSTG [57]: First applies diffusion models to spatiotemporal graph forecasting. By combining STGNNs and diffusion models, it reduces prediction errors and improves uncertainty modeling.
@@
  • TimeGrad [46]: An autoregressive model based on diffusion models. It conducts probabilistic forecasting for multivariate time series and performs well on real-world datasets.
@@
  • CSDI [53]:Utilizes score-based diffusion models for time series imputation. It can leverage the correlations of observed values and also shows remarkable results on prediction tasks.
@@
  • DYffusion [47]: A training method for diffusion models in probabilistic spatiotemporal forecasting. It combines data temporal dynamics with diffusion steps and performs well in complex dynamics forecasting.
@@
  • NPDiff [51]: A general noise prior framework for mobile traffic prediction. It uses the data dynamics to calculate noise prior for the denoising process and achieve effective performance.
C.3 Experimental Configuration

### v5

@@ 4.3 Residual Learning via Diffusion Model
The residual distribution of spatiotemporal data is not independently and identically distributed (i.i.d.) nor does it follow a fixed distribution, such as $\mathcal{N}(0,\mathbf{\sigma})$. Instead, it often exhibits complex spatiotemporal dependence and heterogeneity. We use the diffusion model to focus on learning the distribution of residual $\mathbf{r}^{ta}=\mathbf{x}^{ta}-\mathbb{E}_{\theta}[\mathbf{x}^{ta}|\mathbf{x}^{co}].$ Accordingly, the target data $\mathbf{x}^{ta}$ for diffusion models in Eqs. equation 1, equation 2, and equation 3 is replaced by $\mathbf{r}^{ta}$. We incorporate timestamp information as a condition in the denoising process and concatenate the context data $\mathbf{x}^{co}_{0}$ with noised residual $\mathbf{r}^{ta}_{n}$ as input to capture real-time fluctuations. Notably, no noise is added to $\mathbf{x}^{co}_{0}$ during diffusion training or inference. To model the spatial patterns of the residuals, we propose a scale-aware diffusion process to further distinguish the heterogeneity for different spatial units. In this section, we detail the calculation of $\mathbf{Q}$ and how it is integrated into the scale-aware diffusion process.
@@
(ii) Scale-aware diffusion process. The vanilla diffusion models assume a shared prior distribution $\mathcal{N}(0,I)$ across all spatial locations, failing to capture spatial heterogeneity. To further model such differences, we adopt the technique proposed by Han et al. (2022) to make the residual learning location-specific conditioned on $\mathbf{Q}$}. Specifically, we redefine the noise distribution at the endpoint of the diffusion process as follows:
@@ 4.4 Training and Inference
We adopt a two-stage training procedure: first pretraining a deterministic model to predict the conditional mean, then training a diffusion model to capture the residual distribution (Algorithm 1). During inference, the deterministic model provides the mean prediction, while the diffusion model estimates residuals; their outputs are combined to form the final forecast (Algorithm 2).
@@ 5 Experiments
Baselines. We compare against nine representative state-of-the-art baselines commonly adopted in spatiotemporal modeling, including: Gaussian Process (GP), DeepState (Rangapuram et al., 2018), D3VAE (Li et al., 2022), DiffSTG (Wen et al., 2023), TimeGrad (Rasul et al., 2021), CSDI (Tashiro et al., 2021), DYffusion (Rühling Cachay et al., 2023), TMDM (Li et al., 2024), and NPDiff (Sheng et al., 2025). Detailed descriptions of each baseline are provided in Appendix E.2.
Table 1: Short-term forecasting results in terms of CRPS, QICE, and IS. Bold indicates the best performance, while underlining denotes the second-best. DYffusion is limited to grid-format data, and “-” denotes results that are not applicable.
@@
Metrics. We evaluate performance using two deterministic metrics (MAE, RMSE) and three probabilistic metrics (CRPS, QICE, IS). For QICE, we set $M_{\text{QIs}}=10$ bins following its original design (Han et al., 2022), which offers a balanced trade-off between granularity and stability. For IS, we choose a confidence level of 90% (i.e., $\alpha_{CI}=0.1$) following common practice in spatiotemporal forecasting tasks (Tashiro et al., 2021; Rasul et al., 2021).
Experimental configuration. We define short-term forecasting as predicting the next 12 steps from the previous 12 observations (Sheng et al., 2025; Wen et al., 2023), and long-term forecasting as predicting the next 64 steps from the previous 64 (Yuan et al., 2024; Jin et al., 2023b). As temporal granularity varies across datasets, the actual durations differ. Full configurations are in Appendix E.3.
@@ 5.1 Spatiotemporal Probabilistic Forecasting
Short-term forecasting. Table 1 reports probabilistic metrics, with additional results in Appendix Table 6. CoST consistently outperforms baselines, achieving average improvements of 17.4% in CRPS, 46.6% in QICE, and 16.5% in IS. This indicates superior distribution modeling and more reliable prediction intervals. While not always best on every single metric, CoST remains competitive across datasets. Deterministic results (Table 2, Appendix Table 7) show reductions of 7% in MAE and 6.1% in RMSE, confirming that integration with a strong conditional mean estimator enhances regular pattern capture. In contrast, TMDM underperforms other baselines, mainly because its time-series–oriented design limits generalization to spatiotemporal data, and its focus on long sequences weakens short-term prediction, as evidenced in Appendix Tables 8 and 9. Further experiments on the ETTh1 and ETTh2 time-series datasets show notable performance gains (Appendix Table 11).
Long-term forecasting. As shown in Appendix Table 8, CoST achieves substantial improvements in long-term forecasting under probabilistic metrics, with an improvement of 15.0% and 70.4% in terms of CRPS and QICE. Despite adopting a simple MLP architecture, CoST achieves higher overall accuracy than CSDI, a Transformer-based model tailored for capturing long-range dependencies. Furthermore, it provides significantly better training efficiency and inference speed, as detailed in Section 5.4. In addition, CoST performs well on deterministic metrics (Appendix Table 9), achieving an average reduction of 9.0% in MAE and 11.0% in RMSE compared to the best-performing baseline.
Table 3: Performance of different deterministic backbone models within the CoST framework. “Diffusion (w/o m)” denotes the results obtained using a single diffusion model.
@@
Framework generalization. To demonstrate the generality of CoST, we instantiate it with four representative spatiotemporal forecasting models: STID (Shao et al., 2022), STNorm (Deng et al., 2021), ConvLSTM (Shi et al., 2015), and iTransformer (Liu et al., ). These models cover a diverse set of deep learning architectures, including CNNs, RNNs, MLPs, and Transformers. As shown in Table 3, CoST consistently enhances the performance of these backbones by effectively integrating deterministic and probabilistic modeling. Compared to using a single diffusion model, CoST yields more accurate predictions and better-calibrated uncertainty estimates, validating the framework’s broad applicability and effectiveness.
@@
Case study of SST forecasting. To assess our model’s ability to quantify uncertainty under complex climate dynamics, we evaluate its performance in a key region for ENSO-related Sea Surface Temperature (SST) forecasting. As shown in Figure 3, our model produces high-fidelity SST forecasts that closely match ground truth across both warm pool and cold tongue regions. In addition to accurate mean predictions, it provides well-calibrated uncertainty estimates, revealing elevated variance in the central equatorial Pacific, especially near 0° latitude and 140°–130°W, where sharp thermocline gradients and nonlinear feedbacks make forecasting particularly challenging. These high-uncertainty areas align with known regions of model divergence in climate science (McPhaden et al., 2006; Cane, 2005; Cao et al., 2021), demonstrating that our method delivers both accurate predictions and geophysically consistent uncertainty estimates.
@@ 5.2 Ablation Study
We perform an ablation study to assess the contribution of each proposed module. Specifically, we construct three model variants by progressively removing key components: (w/o s) removes the scale-aware diffusion process; (w/o q) excludes the customized fluctuation scale as a prior; (w/o m) removes the conditional mean predictor, relying solely on the diffusion model. Experiments on two datasets (Appendix Figure 8) show that the deterministic predictor notably improves performance by capturing regular spatiotemporal patterns, while also reducing the diffusion model’s complexity. Adding the customized fluctuation scale further enhances accuracy, indicating its utility in providing valuable fluctuation information across different spatial units. And the scale-aware diffusion process enables the diffusion model to better utilize this condition.
@@ 5.3 Qualitative Analysis
As shown in Figure 5, the ground truth exhibits clear spatiotemporal multi-modality. In Figure 5(a), three peaks likely correspond to different time points or varying states at the same time. CoST accurately captures all three peaks, while CSDI only fits two, showing CoST’s superior multi-modal modeling. In Figure 5(b), both models capture two peaks, but CoST aligns better with the peak spacing in the true distribution, reflecting stronger temporal sensitivity. These strengths arise from CoST’s hybrid design: the diffusion component models residual uncertainty to capture multi-modal traits, while the deterministic backbone learns regular trends. See Appendix E.5.1 for more analysis and results.
@@
Analysis of prediction quality. To intuitively demonstrate the effectiveness of our predictions, we visualize results on the CrowdBJ dataset in Figure 6 (More Cases shown in Appendix Figure 10), comparing our model with the best baseline, CSDI. As shown in Figures 6 (a, c, f), our model, aided by a deterministic backbone, better captures regular spatiotemporal patterns. Meanwhile, the diffusion module enhances uncertainty modeling by focusing on residuals, as reflected in Figures 6 (b, d, e). Beyond sample-level comparison, we evaluate prediction interval calibration via dynamic quantile error curves on CrowdBM and MobileSH (Figure 5). For each confidence level $\alpha$, we compute the corresponding quantile interval and its Prediction Interval Coverage Probability (PICP). Closer alignment with the diagonal (black dashed line) indicates better calibration. Our model consistently outperforms CSDI in this regard.
@@ 5.4 Computational Cost
We benchmark training and inference time (including 50 sampling iterations and pretraining for our mean predictor) on the MobileSH dataset. As shown in Appendix Table 10, CoST is markedly more efficient than existing probabilistic models. The efficiency comes from modeling only the residual distribution with a lightweight denoising network. In contrast, other baselines incur high computational costs by modeling the full data distribution with complex networks, while our approach is well-suited for time-sensitive applications such as mobile traffic prediction.
@@ 6 Conclusion
In this work, we highlight the importance of probabilistic forecasting for complex spatiotemporal systems and propose CoST, a collaborative framework that integrates deterministic and diffusion models. By decomposing data into a conditional mean and residuals, CoST bridges deterministic and probabilistic modeling, accurately capturing regular patterns and uncertainties across diverse spatiotemporal systems. Experiments on ten real-world datasets show that CoST outperforms state-of-the-art methods by 25% on average. Our approach offers an effective solution for combining precise pattern learning with uncertainty modeling in spatiotemporal forecasting.
@@
Reproducibility statement
We provide detailed experimental settings in Section 5 and Appendix E. To facilitate the reproduction of our results, the source code and datasets used in this work have been made publicly available in an anonymous repository at: https://anonymous.4open.science/r/CoST_17116.
Appendix A Usage of LLMs
In compliance with the stated policy, we report the use of a large language model (LLM) as a general-purpose assistance tool in the writing of this paper. Its application was confined to copy-editing and language polishing, such as correcting grammar and syntax. The LLM played no role in the research ideation or the generation of the substantive content of this work.
Appendix B Related Work
B.1 Spatiotemporal deterministic forecasting.
Deterministic forecasting of spatiotemporal systems focuses on point estimation. These models are typically trained with loss functions like MSE or MAE to learn the conditional mean $\mathbb{E}[y|x]$, capturing regular patterns. Common deep learning architectures include MLP-based (Shao et al., 2022; Qin et al., 2023; Zhang et al., 2023), CNN-based (Li et al., 2017; Liu et al., 2018; Zhang et al., 2017), and RNN-based (Bai et al., 2019b; Lin et al., 2020; Wang et al., 2017; 2018) models, valued for their efficiency. GNN-based methods (Bai et al., 2019a; 2020; Geng et al., 2019; Jin et al., 2023a) capture spatial dependencies in graph-based data, while Transformer-based models (Chen et al., 2022; 2021b; Ma et al., 2024; Yu et al., 2020; Boussif et al., 2023) are effective at modeling complex temporal dynamics.
B.2 Spatiotemporal probabilistic forecasting.
The core of probabilistic forecasting lies in modeling uncertainty, aiming to capture the full data distribution (Yang et al., 2024; Tashiro et al., 2021). This is particularly suited for modeling the stochastic nature of spatiotemporal systems. While early methods include classical Bayesian approaches like Gaussian Processes (GP) (Roberts et al., 2013) and influential deep learning models such as DeepAR (Salinas et al., 2020) and DeepStateSpace (Rangapuram et al., 2018), recent advances have explored generative models such as GANs (Jin et al., 2022; Zhang et al., 2021), VAEs (Chen et al., 2021a; De Miguel et al., 2022), and diffusion models (Chai et al., 2024; Lin et al., 2024). Diffusion models, in particular, have gained traction for their ability to model complex distributions with stable training, yielding strong performance in spatiotemporal forecasting (Rasul et al., 2021; Rühling Cachay et al., 2023; Shen & Kwok, 2023; Sheng et al., 2025).
Appendix C Background
C.1 Glossary
@@
NN The number of diffusion steps.
$H$ Height of the grid-based data.
$W$ Width of the grid-based data.
@@
$\mathcal{K}$ $K$ The set of indices for the selected FFT components.
@@ C.2 Spatiotemporal Data
Appendix D Methodology
D.1 Algorithm
@@
Appendix E Experiments
E.1 Datasets
Table 5: The basic information of spatio-temporal data.
@@ E.1 Datasets
Los-Speed Los Angeles Traffic speed 2012/03/01 - 2012/03/07 $207$ 5 min
@@ E.2 Baselines
  • GP (Gaussian Processes):A non-parametric time series forecasting method that models data as a Gaussian process, offering uncertainty estimates and effective modeling of non-linear relationships.
  • DeepState Rangapuram et al. (2018): A deep learning framework for time series forecasting that integrates state space models with neural networks, enabling efficient probabilistic predictions by learning latent states and observation processes.
@@
  • D3VAE (Li et al., 2022): Aims at short-period and noisy time series forecasting. It combines generative modeling with a bidirectional variational auto-encoder, integrating diffusion, denoising, and disentanglement.
@@
  • DiffSTG (Wen et al., 2023): First applies diffusion models to spatiotemporal graph forecasting. By combining STGNNs and diffusion models, it reduces prediction errors and improves uncertainty modeling.
@@
  • TimeGrad (Rasul et al., 2021): An autoregressive model based on diffusion models. It conducts probabilistic forecasting for multivariate time series and performs well on real-world datasets.
@@
  • CSDI (Tashiro et al., 2021):Utilizes score-based diffusion models for time series imputation. It can leverage the correlations of observed values and also shows remarkable results on prediction tasks.
@@
  • DYffusion (Rühling Cachay et al., 2023): A training method for diffusion models in probabilistic spatiotemporal forecasting. It combines data temporal dynamics with diffusion steps and performs well in complex dynamics forecasting.
  • TMDM (Li et al., 2024): TMDM integrates transformers with diffusion models for probabilistic time series forecasting, using transformer-based prior knowledge to enable accurate distribution forecasting and uncertainty estimation.
@@
  • NPDiff (Sheng et al., 2025): A general noise prior framework for mobile traffic prediction. It uses the data dynamics to calculate noise priors for the denoising process and achieve effective performance.
E.3 Experimental Configuration